OpenStudy (miwqueen):

what is the sum of the first 5 terms of a geometric series with a1=10 and r=1/5?

OpenStudy (maryfitzgerald14):

Sn= \frac{a1(1-r^n)}{1-r
} a1=10 r=1/5 n=5 Sn= \frac{10(1-(1/5)^5)}{1-(1/5) } Sn=12.496